OpenStudy (electrochika):

Find the point tangent to the curve at the point defined by the given value of t x = 2cost , y = 2sint , t = pi/4

OpenStudy (asnaseer):

\[y/x=2\sin t/2\cos t=\tan t\]\[y=x*\tan t\]\[dy/dx=\tan t\]so at t=pi/4, we get:\[dy/dx=\tan (\pi/4)=1\]